I don't know where to post this, hope someone can help
p=c+b+4a solve for c

Found 2 solutions by richwmiller, mananth:
Answer by richwmiller(17219)

subtract everything but the c that is one the right side so that only c is on the right

Answer by mananth(16949)

p=c+b+4a solve for c
since you have to solve for c you have to keep c on one side of the equation and move the other terms to the other side of c.
to remove b and 4a from the right side add -b and -4a on the right side and the left side of the equation, so that the equation does not change.
p-b-4a = c+b+4a -b-4a
p-b-4a=c
Ananth